Please help!
I have to solve these equations by elimination:
#1 2x + 3y = 12 and y = 4/3x -2
#2 y = 1/2 x + 2 and y = 2x -1

Answer by jennywecas(28)

#1 2x + 3y = 12 and y = 4/3x -2
first multiply (y= 4/3x-2)by -3
you get -3y=-4x+6
2x=-4x+18
6x=18
x=3
6+3y=12
3y=6
y=2
ANSWER IS (3,2)
#2 y = 1/2 x + 2 and y = 2x -1
y-2=1/2x
-y-1=-2x
-3=-1.5x
x=2
y=2*2-1
y=3
ANSWER IS (2,3)
x=2